OpenStudy (anonymous):

The zero of F-1(x) in F(x) = x + 3 is A.-3 B.0 C.3

OpenStudy (aravindg):

first find f-1(x)

OpenStudy (aravindg):

to get f-1(x) put y=f(x)=x+3 then write x in terms of y